# EdgeOL: Efficient in-situ Online Learning on Edge Devices

### 30 Jan 2024 v1

@@ 4.3. Handling Multiple Scenario Changes
4.4. Utilizing Unlabeled Data
It is possible that some streaming-in training data is not labeled. To fine-tune the model in this case, EdgeOL applies a semi-supervised learning technique (chapelle2009semi, ). In each fine-tuning round, EdgeOL first fine-tunes the model using unlabeled data via the self-supervised learning (chen2020simple, ) to improve the feature extraction ability of the model. Then, EdgeOL fine-tunes the model by supervised learning using the labeled data to improve the model’s performance in the target task (e.g., image classification for a particular dataset).
@@ 5. Evaluation
In this section, we evaluate the proposed EdgeOL framework using popular online learning workloads from both CV and NLP domains. The evaluation is conducted under two scenario change types: task change and input distortion.
In CV domain, we employ two popular CNN models ResNet-50 (Res50 for short) and MobileNetV2 (MBV2 for short), as well as a vision transformer model DeiT (tiny version) (pmlr-v139-touvron21a, ). To evaluate EdgeOL on task change, we use two test cases “ImgNet to CF10” and “ImgNet to CalT101”. The models were originally well-trained on the ImageNet dataset and then targeted to serve CIFAR-10/Caltech101 dataset with the corresponding streaming-in training data and inference requests. To evaluate EdgeOL on input distortion, we use the ImageNet-C/CIFAR-10-C dataset (hendrycks2019benchmarking, ), which has the same image classes as ImageNet/CIFAR-10, but its image involves distortion such as fog and snow. Two test cases “ImgNet to ImgNet-C” and “CF10 to CF10-C” are used in our evaluation. We also evaluate EdgeOL on CORe50 benchmarks (lomonaco2020rehearsal, ) for multiple scenario change cases in Section 5.4, which include both task changes and input distortions. In NLP, we evaluate EdgeOL by fine-tuning the BERT-base model (kenton2019bert, ) (originally well-trained on English Wikipedia and the Brown Corpus) using two popular datasets, MRPC (dolan2005automatically, ) and CoLA (warstadt2019neural, ).
Each dataset contains training and testing sets, and a portion of training data (5%) is randomly separated to form a validation dataset as discussed in Section 4.1.2. The experimental setup is the same as in Section 3.1. We assume a total of 500 inference requests, and we provide a sensitivity study on different numbers of inference requests in Section 5.3. All reported accuracies are the average over 5 runs using different random seeds. Unless otherwise stated, the accuracy results refer to the average inference accuracy.
@@
We use immediate online learning as our baseline, where the models are fine-tuned once training data is available. To showcase the benefits of fine-tuning, we present the average inference accuracy achieved by using the pre-trained model without fine-tuning (denoted as “w/o FT”) to serve all the inference requests. We also compare EdgeOL with state-of-the-art efficient training methods (Section 5.2), including layer freezing methods i) Egeria (wang2023egeria, ) and ii) AutoFreeze (liu2021autofreeze, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ).
5.1. Main Results
5.1.1. CV Tasks
Figure 1112, and Table 3 show the overall execution time, energy consumption, and average inference accuracy of the immediate online learning, model without fine-tuning (accuracy only), and our proposed frameworks in CV domain. The execution time and energy consumption are normalized to $Immed.$ As shown in Table 3, fine-tuning can uniformly outperform the model without fine-tuning in accuracy, as it can ensure adaptiveness for the model.
@@ 5.1.1. CV Tasks
Dataset
ImgNet to ImgNet to ImgNet to CF10 to
CF10 CalT101 ImgNet-C CF10-C
@@
87.79
73.27
71.77
93.82
@@
Table 4. Computation cost of entire online learning process. Test case:“ImgNet to CF10”. Method Computation (TFLOPs) Res50 MBV2 Immed. 1,236 96 EdgeOL 767 58
[Uncaptioned image] Figure 13. Computation cost for 1 iteration and memory usage as fine-tuning proceeds.
[Uncaptioned image] Figure 14. Convergence of ResNet50 model on “ImgNet to CF10” test case.
@@
DAF. As shown in Figures 11 and 12, DAF saves average 74%, peak 86% execution time, and average 54%, peak 77% energy compared to $Immed$. These savings come from merging and delaying certain fine-tuning rounds, which can effectively reduce the execution overheads (by 95% on average), including system initialization and model loading & saving (orange bar and blue bar in Figure 11). As shown in Table 3, despite the impressive gains in time and energy efficiency, DAF only incurs a minor 0.16% accuracy drop compared to $Immed$. This is because it adaptively determines the fine-tuning frequency that fits the current scenario.
@@
SimFreeze. SimFreeze reduces average 8%, peak 15% execution time and saves average 20%, peak 33% energy compared to $Immed.$, as shown in Figures 11 and 12. These gains stem from the 40% average savings in model computation (i.e., forward and backward propagation) through layer freezing. Notably, SimFreeze also delivers significantly higher accuracy, a 1.91% average increase over $Immed.$, as shown in Table 3. The reasons are two-fold: First, SimFreeze accelerates model convergence (shown in Figure 14) as freezing layers reduce the number of model weights being trained. Second, SimFreeze avoids excessive adaptation to training data by freezing well-trained layers.
@@
EdgeOL. EdgeOL combines DAF and SimFreeze. From Figures 11, 12, and Table 3, compared to $Immed.$, EdgeOL saves average 82%, peak 91% execution time and average 74%, peak 89% energy, and improves accuracy by an average of 1.70%. Note that EdgeOL shows more time (average 4%) and energy (average 10%) savings in the input distortion test cases, as their scenario changes are less pronounced, allowing greater optimization potential in both inter- and intra-tuning.
@@
Computation Cost and Memory Usage. Table 5.1.1 shows the computation cost reduction. Figure 13 (left) shows the computing cost per iteration as the fine-tuning proceeds. Note, computation cost reduction comes from SimFreeze, as DAF only delays and merges fine-tuning rounds. EdgeOL also saves memory since freezing layers can reduce the intermediate data generated during the computation. As shown in Figure 13 (right), EdgeOL can reduce the memory usage by 40% for ResNet50 and MobileNetV2.
@@
Model Convergence Speed. Figure 14 plots the model convergence during entire online learning. We observe that our EdgeOL helps the model converge faster, leading to a higher accuracy compared to immediate online learning.
@@
Overheads. The major overhead of EdgeOL is the CKA calculation in SimFreeze. This overhead is introduced by i) a forward propagation using a batch of data to get the output feature maps, ii) CKA calculation for active layers using the obtained output feature maps. Fortunately, many layers will be frozen as training proceeds, so the computation of CKA decreases over time. Specifically, in our evaluation, SimFreeze incurs $<$3% additional energy for CKA computation, which is negligible compared to 74% energy benefit from EdgeOL. The reported results include these overheads.
5.1.2. NLP Tasks
We evaluate the EdgeOL framework on NLP tasks to showcase its generalizability. As shown in Table 5, EdgeOL achieves over 30% higher accuracy compared to the model without fine-tuning. When compared to immediate online learning, EdgeOL offers an average reduction of 76% in execution time and 63% in energy consumption, while increasing the accuracy by 1.76% on average. These results demonstrate the generalizability and superiority of EdgeOL.
@@ 5.1.2. NLP Tasks
MRPC CoLA
Acc.
Time
Energy
Acc.
Time
Energy
w/o FT
31.61
N/A
N/A
30.87
N/A
N/A
@@
74.38
873
4.11
51.05
2,278
9.26
Figure 15. Results under the different number of inference requests.
Figure 16. Results under different arrival distributions.
Figure 15. Results under the different number of inference requests.
Figure 16. Results under different arrival distributions.
Figure 17. Results under different CKA variation threshold.
5.2. Comparison with State-of-the-art Efficient Learning Methods
We compare EdgeOL with state-of-the-art efficient training methods, including layer freezing methods i) Egeria (wang2023egeria, ) and ii) AutoFreeze (liu2021autofreeze, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ). Results are presented in Table 6. All these methods do not consider optimizations of inter-tuning, which significantly limits their benefits in execution time, energy consumption, and accuracy. For a thorough comparison, we integrate our inter-tuning optimization, DAF, into all methods with identical configurations. Table 6 shows that even with DAF integration, EdgeOL still consistently outperforms all these methods. As shown in the table, EdgeOL provides 2.1$\times$, 2.2$\times$, 2.6$\times$, and 2.2$\times$ energy savings, respectively, while delivering 1.76%, 2.10%, 2.17%, and 1.89% higher accuracy.
EdgeOL outperforms Egeria and AutoFreeze due to its more flexible and finer-grained layer-freezing approach. Spe- cifically, EdgeOL assesses layers individually rather than in modules (i.e., layer blocks), and it freezes all identified converged layers without forcing layers to be frozen sequentially from front to back. Hence, it avoids overtraining already converged layers in the middle of a non-converged module or after a non-converged layer. In contrast to RigL, EdgeOL effectively addresses sparse training challenges, such as GPU underutilization and workload imbalance. Compared to Ekya, EdgeOL eliminates the inefficiency of Ekya’s trial-and-error method in training configuration (e.g., which layers to freeze), ensuring more effective and efficient performance improvements.
@@ 5.2. Comparison with State-of-the-art Efficient Learning Methods
ImgNet to CF10 ImgNet to ImgNet-C
Acc.
Time
Energy
Acc.
Time
Energy
@@
86.22
4,130
16.16
69.67
8,692
34.27
EdgeOL 87.79 3,461 13.56 71.77 6,877 25.29
@@ 5.3. Sensitivity Analysis
Number of inference requests. Figure 17 shows the average inference accuracy and energy consumption under different numbers of inference requests. Note that, all the inference requests are arriving following a Poisson distribution (mattson2020mlperf, ). All results in this section are obtained by fine-tuning ResNet50 on “ImgNet to CF10”. EdgeOL consistently achieves higher accuracy than $Immed.$, while consuming significantly less energy. The figure also reveals that the energy saving offered by EdgeOL increases as the total number of inference requests decreases. This occurs because when the number of inference requests decreases, EdgeOL (achieved by DAF) will decrease fine-tuning frequency, thereby reducing the energy from execution overheads such as system initialization, as explained in Section 3.2.
Data Arrival distribution. In addition to the Poisson distribution, we also evaluate EdgeOL under different arrival distributions for both training data and inference requests, including the uniform distribution (kuipers2012uniform, ), normal distribution (altman1995statistics, ), and a real-world trace from Video Timeline Tags dataset (huang2020multimodal, ). As depicted in Figure 17, EdgeOL consistently excels in both accuracy and energy consumption compared to $Immed$., showing that EdgeOL is adept at handling a variety of scenarios with different data arrival distributions.
@@
5.4. Handling Multiple Scenario Changes
We evaluate EdgeOL when there are multiple model deployment scenario changes throughout online learning. We employ two CORe50 benchmarks NICv2_79 and NICv2_391 (lomonaco2020rehearsal, ), which are widely used in testing the model’s adaptability in dynamically changing scenarios. “NIC” stands for New Instance and Classes, and the numbers (79 and 391) indicate the total scenarios in each benchmark. The scenarios appear one after one, each of which introduces both i) new classes of data and ii) instances of existing classes but with new patterns (e.g., different environmental conditions such as changes in illumination and background).
In Table 7, we compare EdgeOL with layer freezing method Egeria (wang2023egeria, ) and efficient online learning framework Ekya (bhardwaj2022ekya, ). Following the experimental setup described in Section 5.2, both Egeria and Ekya utilize our inter-tuning level optimization, DAF, to ensure a fair comparison. EdgeOL achieves energy savings of 2.3$\times$ and 2.0$\times$ over Egeria and Ekya, respectively, while also providing 1.15% and 1.00% higher accuracy. The superior performance of EdgeOL in handling multiple scenario changes can be attributed to the Reset & Resume design (Section 4.3). When encountering a new scenario, it enables a quick adaptation by increasing the fine-tuning frequency and selectively resuming fine-tuning only for those frozen layers that become unstable in the new scenario.
Table 7. Experiments on handling multiple scenario changes.
@@ 5.4. Handling Multiple Scenario Changes
NICv2_79 NICv2_391
Acc.
Time
Energy
Acc.
Time
Energy
Res50
DAF (base)
56.94
9,718
34.07
54.20
10,113
34.40
Egeria (wang2023egeria, )
56.60
8,634
29.19
54.23
9,334
29.62
Ekya (bhardwaj2022ekya, )
57.01
8,586
28.80
54.54
9,281
29.00
EdgeOL 57.86 7,482 23.50 55.73 7,720 24.08
MBV2
DAF (base)
53.28
3,066
10.00
50.87
3,224
10.21
Egeria (wang2023egeria, )
53.60
2,796
8.55
50.95
2,916
8.56
Ekya (bhardwaj2022ekya, )
52.78
2,759
8.31
51.22
2,903
8.37
EdgeOL 54.11 2,483 6.70 51.94 2,633 6.81
DeiT
DAF (base)
56.42
9,845
37.19
53.63
10,583
39.91
Egeria (wang2023egeria, )
56.48
8,302
33.90
53.11
8,040
33.02
Ekya (bhardwaj2022ekya, )
56.73
7,885
31.77
53.56
8,844
34.18
EdgeOL 57.51 6,490 25.66 54.70 7,090 28.33
Table 8. Experimental results in semi-supervised learning. The test case is “ImgNet to CF10”.
Model
Method
Acc.
Time (Sec.)

### 16 Mar 2024 v2

@@ 5. Evaluation
In this section, we evaluate the proposed EdgeOL framework using popular online learning workloads from both CV and NLP domains.
5.1. Experimental Setup
Platform: We use the NVIDIA Jetson Xavier NX as our platform and choose the 15W 6-Core power mode with maximum GPU speed.
Model and dataset: In the CV domain, we employ two popular CNN models ResNet50 (Res50 for short) and MobileNetV2 (MBV2 for short), as well as a vision transformer model DeiT (tiny version) (pmlr-v139-touvron21a, ). We employ three benchmarks NC, NICv2_79, and NICv2_391 (lomonaco2020rehearsal, ) from the CORe50 dataset for evaluation, which contain 9, 79, and 391 scenarios, respectively. CORe50 is a popular dataset that is widely used in several prior continuous online learning works (logacjov2021learning, ; mandivarapu2020self, ; graffieti2022continual, ; pellegrini2021continual, ). In the NC benchmark, each scenario introduces new classes of data on top of existing classes. On the other hand, in NIC benchmarks, each scenario can introduce either i) new classes of data, ii) instances of existing classes but with new patterns (e.g., different environmental conditions such as changes in illumination and background), or iii) a combination of both. We also use another widely-used benchmark S-CIFAR-10 (buzzega2020dark, ; wang2022sparcl, ) to evaluate EdgeOL, where the CIFAR-10 (krizhevsky2009learning, ) dataset is split into 5 scenarios, each consisting of 2 distinct data classes. In the NLP domain, we employ the BERT-base model (kenton2019bert, ) and the 20News benchmark used in several prior works (kirkpatrick2017overcoming, ; ke2021adapting, ; ke2021achieving, ; sun2019lamol, ), where the 20News dataset (lang1995newsweeder, ) is split into 10 scenarios, each containing 2 data classes.
Fine-tuning Setting: In our experiments, the model is originally well-trained in the first scenario. In the subsequent scenarios that appear sequentially, it will be online fine-tuned with corresponding training data and meanwhile serve inference requests. The CopyWeights with Re-init (CWR) technique proposed by the CORe50 benchmark paper is by default applied in the experiments to mitigate the catastrophic forgetting problem (zenke2017continual, ; li2019learn, ).
In each scenario during the entire online learning process, both the training data and inference requests arrive continuously over time. The arrival granularity of training data is 1 batch each time and the training batch size is fixed to 16 to avoid out-of-memory errors. We assume a total of 500 inference requests across all scenarios. The arrival rate for both the training data and inference requests follows a Poisson distribution to mimic real application scenarios (mattson2020mlperf, ). We also provide a sensitivity study on different numbers of inference requests and different arrival distributions in Section 5.4. Each dataset contains training and testing sets, and a portion of training data (5%) is randomly separated to form a validation dataset as discussed in Section 4.1.2.
@@ 5.1. Experimental Setup
Baseline and SOTA Comparisons: We use immediate online learning as our baseline, where the models are fine-tuned once training data is available. It provides the highest accuracy over all the online learning strategies with fixed fine-tuning frequency (see Figure 4 for details). We also compare EdgeOL with state-of-the-art efficient training methods (Section 5.3), including layer freezing methods i) Egeria (wang2023egeria, ) and ii) SlimFit (ardakani2023slimfit, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ).
Metrics: We use three metrics for evaluation: overall fine-tuning execution time, overall energy consumption, and average inference accuracy. The overall fine-tuning execution time and energy consumption refer to the total time and energy costs of all scenarios during the entire online learning process. They sum up the fine-tuning execution time and energy consumption of all fine-tuning rounds. The average inference accuracy is the average of accuracies over all inference requests in all scenarios. All reported results are the average over 5 runs using different random seeds. Unless otherwise stated, the accuracy results refer to the average inference accuracy.
@@
Benchmark
NC
NICv2_79
NICv2_391
S-CIFAR-10
@@
70.69
62.95
52.77
85.80
5.2. Main Results
5.2.1. CV Tasks
Figures 1212, and Table 3 show the overall execution time, energy consumption, and average inference accuracy of the immediate online learning and our proposed frameworks in CV domain. The execution time and energy consumption are normalized to $Immed.$
@@ 5.2.1. CV Tasks
Table 4. Computation cost of entire online learning process of the NC benchmark. Method Computation (TFLOPs) Res50 MBV2 Immed. 4,746 367 EdgeOL 3,037 124
[Uncaptioned image] Figure 13. Memory usage at the beginning and the end of online learning.
[Uncaptioned image] Figure 14. Convergence of ResNet50 in one of the scenarios in NC benchmark.
@@
DAF. As shown in Figures 12 and 12, DAF saves average 50%, peak 68% execution time, and average 31%, peak 62% energy compared to $Immed$. These savings come from merging and delaying certain fine-tuning rounds, which can effectively reduce the execution overheads (by 92% on average), including model loading, saving, and system initialization (e.g., model compilation). As shown in Table 3, despite the impressive gains in time and energy efficiency, DAF only incurs a minor 0.22% accuracy drop compared to $Immed$. This is because it adaptively determines the fine-tuning frequency that fits the current situation.
@@
SimFreeze. SimFreeze reduces average 15%, peak 23% execution time and saves average 22%, peak 26% energy compared to $Immed.$, as shown in Figures 12 and 12. These gains stem from the 35% average savings in model computation (i.e., forward and backward propagation) through layer freezing. Notably, SimFreeze also delivers significantly higher accuracy, a 1.96% average increase over $Immed.$, as shown in Table 3. The reasons are two-fold: First, SimFreeze accelerates model convergence (shown in Figure 14) as freezing layers reduce the number of model weights being trained. Second, SimFreeze avoids excessive adaptation to training data by freezing well-trained layers.
@@
EdgeOL. EdgeOL combines DAF and SimFreeze. From Figures 12, 12, and Table 3, compared to $Immed.$, EdgeOL saves average 64%, peak 78% execution time and average 52%, peak 73% energy, and improves accuracy by an average of 1.75%. Note that EdgeOL shows more time and energy savings in NC and S-CIFAR-10 benchmarks, as their scenario changes are less frequent (8 and 4 vs. 78 and 390), allowing greater optimization potential in both inter- and intra-tuning.
@@
Computation Cost and Memory Usage. Table 5.2.1 shows the computation cost reduction. Note that computation cost reduction comes from SimFreeze, as DAF only delays and merges fine-tuning rounds. EdgeOL also saves memory since freezing layers can reduce the intermediate data generated during the computation. As shown in Figure 13, EdgeOL can reduce the memory usage by 40% for ResNet50 and MobileNetV2.
@@
Model Convergence Speed. Figure 14 plots the model convergence during online learning in one of the scenarios. We observe that our EdgeOL helps the model converge faster as layer freezing effectively reduce the number of model weights being trained, leading to a higher accuracy compared to immediate online learning.
@@
Overheads. The major overhead of EdgeOL is the CKA calculation in SimFreeze. This overhead is introduced by i) a forward propagation using a batch of data to get the output feature maps, ii) CKA calculation for active layers using the obtained output feature maps. Fortunately, many layers will be frozen as training proceeds, so the computation of CKA decreases over time. Specifically, in our evaluation, SimFreeze incurs $<$2% additional energy for CKA computation, a minor amount when compared to 52% energy benefit from EdgeOL. All the reported results have included all the overhead.
@@
Accuracy (%)
Time (minute)
Energy (Wh)
@@
66.95
110
30.02
5.2.2. NLP Tasks
We further evaluate the EdgeOL framework on NLP tasks to showcase its generalizability. As shown in Table 5, when compared to the immediate online learning approach, EdgeOL offers a reduction of 67% in execution time and 54% in energy consumption, while increasing the accuracy by 1.52%. These results demonstrate the generalizability and superiority of EdgeOL.
@@ 5.2.2. NLP Tasks
NC NICv2_391
Acc. (%)
Energy (Wh)
Acc. (%)
Energy (Wh)
@@
68.96
61.04
51.06
75.51
EdgeOL 70.69 49.66 52.77 65.74
5.3. Comparison with State-of-the-art Efficient Learning Methods
We compare EdgeOL with state-of-the-art efficient training methods, including layer freezing methods i) Egeria (wang2023egeria, ) and ii) SlimFit (liu2021autofreeze, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ). Results are presented in Table 6. Since all these methods do not consider optimizations of inter-tuning, which significantly limits their benefits in efficiency and accuracy. For a thorough comparison, we integrate our inter-tuning optimization, DAF, into all methods with identical configurations. Table 6 shows that even with DAF integration, EdgeOL still consistently outperforms all these methods, providing 2.1$\times$, 2.2$\times$, 2.8$\times$, and 2.0$\times$ energy savings, respectively, while delivering 1.78%, 2.18%, 2.33%, and 1.50% higher accuracy.
EdgeOL outperforms Egeria due to its more flexible and finer-grained layer-freezing approach. Specifically, EdgeOL assesses layers individually rather than in modules (i.e., layer blocks), and it freezes all identified converged layers without forcing layers to be frozen sequentially from front to back. Hence, it avoids overtraining already converged layers in the middle of a non-converged module or after a non-converged layer. Against SlimFit, EdgeOL’s advantage lies in the use of a more reliable metric: layer representational similarity. This directly analyzes layer outputs, offering a more accurate assessment than indirect methods like monitoring weight update magnitudes, which SlimFit employs. In contrast to RigL, EdgeOL effectively addresses sparse training challenges, such as GPU underutilization and workload imbalance. Compared to Ekya, EdgeOL eliminates the inefficiency of Ekya’s trial-and-error method in training configuration (e.g., which layers to freeze), ensuring more effective and efficient performance improvements.
Figure 15. Results under different number of inference requests.
Figure 16. Results under different arrival distributions.
Figure 17. Results under different CKA variation threshold.
@@ 5.4. Sensitivity Analysis
Number of inference requests. Figure 15 shows the average inference accuracy and energy consumption under different numbers of inference requests. Note that, all the inference requests are arriving following a Poisson distribution (mattson2020mlperf, ). All results in this section are obtained on ResNet50 and NC benchmark. EdgeOL consistently achieves higher accuracy than $Immed.$, while consuming significantly less energy. The figure also reveals that the energy saving offered by EdgeOL increases as the total number of inference requests decreases. This occurs because when the number of inference requests decreases, EdgeOL (achieved by DAF) will decrease the fine-tuning frequency, thereby reducing the energy from execution overheads such as system initialization, as explained in Section 3.1.
Data Arrival distribution. In addition to the Poisson distribution, we also evaluate EdgeOL under different arrival distributions for both training data and inference requests, including the uniform distribution (kuipers2012uniform, ), normal distribution (altman1995statistics, ), and a real-world trace from Video Timeline Tags dataset (huang2020multimodal, ). As depicted in Figure 16, EdgeOL consistently excels in both accuracy and energy consumption compared to $Immed$., showing that EdgeOL is adept at handling a variety of situations with different data arrival distributions.
@@
Table 7. Experimental results (NC benchmark) in semi-supervised learning.
@@
Accuracy (%)